OpenStudy (anonymous):

1) pro basket ball player has a vertical leap of 30 inches, what is the hang time ? The hang time function is v=48T^2. hANG TIME IS------- in seconds. can i get a conformation on my answer for this problem? I need to know what the hang time is in seconds. I came up with 0.79 seconds but am in doubt, need reaffirming. thank you!

OpenStudy (anonymous):

did you solve \[30=48t^2\] \[t^2=\frac{30}{48}=\frac{5}{8}\] \[t=\sqrt{\frac{5}{8}}\]

OpenStudy (anonymous):

then a calculator. i will check

OpenStudy (anonymous):

you got it!

OpenStudy (anonymous):

thanks again satalite

OpenStudy (anonymous):

yw